# QiMeng-SALV: Signal-Aware Learning for Verilog Code Generation

### v1

@@ 1 Introduction
In contrast to high-level program language that defines sequential execution behavior, Verilog code specifies the structural interconnection of hardware gates and wires, granting inherent independence between different output signals. This characteristic enables scenarios where individual signal implementations remain functionally correct despite errors in the overall module implementation.
As shown in Fig.1, the faulty implementation of the signal a results in total malfunction of the top module, whereas the signal d maintains proper functionality. Due to Verilog’s inherent signal independence and parallel processing nature, we can easily extract a certain signal implementation from the top module through abstract syntax tree (AST). Such partially valid signal-level implementations can be leveraged within RL frameworks to derive meaningful functional rewards, thereby expanding the effective sample space for RL optimization.
Based on the above analysis, in this work, we propose Signal-Aware Learning for Verilog code generation (QiMeng-SALV) by leveraging code segments of functionally correct output signal to optimize RL training. The key insight of QiMeng-SALV is to extract verified signal-aware implementations in partially incorrect modules, so as to enhance the extraction of meaningful functional rewards. Specifically, to verify the functional correctness of signals, some random test inputs are generated to compare signal-level outputs between generated module and reference module in the training data.
Additionally, abstract syntax tree (AST) analysis is employed to precisely identify preferred and dispreferred code segments, thus these partially correct signal-level implementation can be utilized to generate meaningful functional rewards.
Finally, to enable the model to learn correct signal implementations from erroneous modules, we introduce signal-aware DPO which is based on DPO, a method renowned for its computational efficiency, rapid convergence, and strong performance.
Unlike standard DPO that computes probabilities for all module code tokens, signal-aware DPO exclusively calculates token probabilities for the correct signal-level code segments from the preferred and dispreferred samples, thereby preventing noise and interference from incorrect signals during RL training. Crucially, the proposed QiMeng-SALV expands the effective training dataset by incorporating all modules containing any valid signal implementations, regardless of overall module correctness. This represents a fundamental shift from conventional module-level to fine-grained signal-level optimization in RL-based Verilog code generation, addressing the issue of insufficient functional rewards in RL training. To the best of our knowledge, this is the first fine-grained signal-level reinforcement learning algorithm designed for Verilog code generation.
@@ 2.2 RL Training for LLM
In the domain of Verilog generation, current RL-based training approaches face two key challenges: (1) highly rely on structural similarity metrics wang2024veriseek ; wu2024itertl that cannot properly evaluate functional equivalence between different implementations, and (2) the scarcity of high-quality SFT data limits the initial model capability needed for effective RL optimization. While recent work has explored various reward formulations chen2021evaluatingeg2 ; le2022coderl ; wang2024veriseek and feedback mechanisms liu2023rltf ; dou2024stepcoder ; cui2024origen , these approaches often introduce evaluation biases, limiting their reliability in assessing functional correctness.
VeriPrefer wang2025insights mitigates such biases by employing module-level functional rewards during reinforcement learning. Our approach differs fundamentally: we adopt signal-level functional rewards, allowing partially correct modules to still provide valuable feedback during training when they contain correctly implemented signals. Consequently, our method delivers denser and more informative reward signals, leading to superior reinforcement learning performance.
@@ 3.2 Signal-aware Verification
First, we use Yosys wolf2013yosys to analyze the reference module head, automatically generating $N$ sets of input signals at equal time intervals.
Second, the input signals are simultaneously applied to the reference module $y$ and generated modules $E=\{y_{1},y_{2},...,y_{k}\}$. By comparing the output signals produced by the generated modules with the corresponding output signals from the reference module, we can verify the correctness of generated modules, and obtain the pairs $\{(y_{1},c_{1}),(y_{2},c_{2}),..,(y_{k},c_{k})\}$, where $c_{i}$ is the correct output signals for generated module $y_{i}$. Notably, only output signals exhibiting complete matching across all $N$ sets of input signals are verified as correct, while any discrepancies result in classification as erroneous signals.
Third, to select preference samples suitable for DPO training, we choose sample pairs that there exists a set of contrast signals that is correct in the preferred sample but incorrect in the dispreferred sample. This can be formally expressed as:
@@ 3.3 Signal-aware Code Extraction
Based on the above analysis, in the signal-aware code extraction stage, for given target signals, we parse the Abstract Syntax Tree (AST) to establish dependency relationships between signals, and then extract the complete code segment corresponding to the target signals, as depicted in Fig.2 (c). Specifically, given the module code of the target signal, we use Yosys wolf2013yosys to parse the module code and obtain the AST. Subsequently, we analyze the AST to derive the dependency relationships between all the output signals and intermediate signals defined in the module code, forming a signal topology graph. By analyzing this topology graph and performing backward traversal from the target output signal leaf nodes, we obtain the dependent signals of the target signal. Based on the locations of these dependent signals in the AST, we retain their corresponding code segments, thereby obtaining the complete code segment of the target signal.
To facilitate the model’s learning by using code segment of contrast signal during DPO training, given the contrast signals $c$, we isolates the relevant code segment $S_{w}^{c}$ from the preferred sample $y_{w}$. Meanwhile, to enable comparison with the preferred sample, we also extract the code segment $S_{l}^{c}$ of contrast signal $c$ from the dispreferred sample $y_{l}$.
@@ 3.4 Signal-aware DPO Training
where $y_{w}$ and $y_{l}$ represents the preferred and dispreferred sample respectively. $S_{w}^{c}$ denotes code segments related to contrast signals $c$ in preferred samples, while $S_{l}^{c}$ represents the same signal related segments in dispreferred samples. $\pi_{\theta}$ is the policy model to be optimized and $\pi_{\text{ref}}$ is the reference model used for regularizing $\pi_{\theta}$ with Kullback-Leibler divergence and $\beta$ is a constant to control the degree of regularization.
Through the objective function in Eq. (2), signal-aware DPO shift DPO’s learning focus from entire modules to individual signals, refining the learning granularity and increasing the quantity of functional correctness rewards in DPO, thereby enhancing optimization effectiveness.
@@ 4.1 Experiment Setup
Datasets. Our training dataset is sourced from CodeV zhao2024codev , consisting of 165K samples obtained by crawling all publicly available Verilog module code on GitHub. However, we observed that a subset of these modules are not syntactically valid, prompting us to filter out such cases and retain a cleaned dataset of 135K samples.
The 135K dataset is initially employed to fine-tune the general-purpose code generation model Qwen2.5 Coder Instruct hui2024qwen2 , yielding a naive code generator. For every prompt in the 135K dataset, this generator produces 5 candidate module codes, while the corresponding module codes in the dataset serve as reference codes to validate signal correctness during Signal-aware Verification.
Training Settings. In training, we perform full-parameter fine-tuning for 2 epochs during the SFT, followed by about 1 epoch (7000 steps) of LoRA-basedhu2022lora fine-tuning in the Signal-aware DPO. Optimization is carried out using the Adam kingma2014adam optimizer with a cosine annealing learning rate schedule, where the initial learning rates are set to 1.0e-5 for SFT and 5.0e-6 for Signal-aware DPO. The training configuration employs a global batch size of 64 and a maximum sequence length of 2048 tokens.
@@
Baseline Methods In our experimental evaluation, we conduct a comprehensive comparison between our proposed method and several baseline approaches, categorized into three groups:
1) General-purpose Foundation Models: GPT-3.5, GPT-4o, GPT-4 achiam2023gpt4 , and DeepSeek-v3 liu2024deepseekv3 , which demonstrate broad capabilities across diverse domains. 2) General Code Models: CodeQwen1.5 qwen1.5 , Qwen2.5 Coder Instruct hui2024qwen2 and Deepseek Coder guo2024deepseekcoder , possessing strong coding proficiency but lacking Verilog-specific optimization. 3) Domain-Specialized Verilog Models: Including RTL Coder liu2024rtlcoder , CodeV zhao2024codev , Origen cui2024origen , VeriSeek wang2024veriseek and VeriPrefer wang2025insights which employ module-level rewards in reinforcement learning.
@@ 4.3 Ablation Study
In this section, we present a systematic ablation study to evaluate our proposed QiMeng-SALV. Given that RTLLM contains more complex problems than VerilogEval, making it more suitable for revealing nuanced performance differences among models, we focus our ablation analysis on the RTLLM benchmark. Our investigation spans two key dimensions: 1) configuration variations of Signal-aware DPO training, and 2) different training stage implementations of QiMeng-SALV.
Ablation of configuration variations of signal-aware DPO Unlike standard DPO, which is limited to processing complete correct modules and must consider all response tokens, our signal-aware DPO can not only handle complete correct modules but also extract functional reward in partial correct modules containing correct signal implementations. Therefore, we design three different options to evaluate the efficiency of our method: 1) Complete correct datasets (263k): all of preferred sample in DPO training data are complete correct modules. 2). Partial correct datasets (110k): all of preferred sample in DPO training data are partial correct modules containing both correct and incorrect signals. 3). Filter incorrect signals: determines whether erroneous signals in preferred samples participate in DPO computation. As evidenced by the results in Table 3, standard DPO trained exclusively on complete correct datasets achieves superior performance compared to training on partially correct datasets. Interestingly, combining both complete and partially correct datasets during training leads to degraded performance relative to using only complete correct datasets. This degradation likely stems from the presence of incorrect signal implementations in partially correct module, which introduce noise and hinder the model’s ability to discern meaningful reward signals. By contrast, our proposed signal-aware DPO method, which actively filters out incorrect signal implementations, substantially improves performance in both scenarios: training on partially correct datasets alone and training on mixed datasets. These findings highlight two key advantages of signal-aware DPO: (1) its ability to leverage a broader spectrum of training data, and (2) its robustness against noisy signal implementations through effective filtering.
@@
Scaling on Signal-aware DPO training datasets
To examine the influence of training dataset size on Signal-aware DPO’s efficacy, we conducted experiments by training models with preference datasets of different scales. Figure 3 (b) reveals a consistent upward trend in pass@1, pass@5, and pass@10 metrics as the dataset size grows. These results indicate that incorporating more meaningful functional rewards during reinforcement learning training significantly contributes to improved model performance.

### v2

@@ 1 Introduction
In contrast to high-level program language that defines sequential execution behavior, Verilog code specifies the structural interconnection of hardware gates and wires, granting inherent independence between different output signals. This characteristic enables scenarios where individual signal implementations remain functionally correct despite errors in the overall module implementation. As shown in Fig.1, the faulty implementation of the signal a results in total malfunction of the top module, whereas the signal d maintains proper functionality. Due to Verilog’s inherent signal independence and parallel processing nature, we can easily extract a certain signal implementation from the top module through abstract syntax tree (AST). Such partially valid signal-level implementations can be leveraged within RL frameworks to derive meaningful functional rewards, thereby expanding the effective sample space for RL optimization.
Based on the above analysis, in this work, we propose Signal-Aware Learning for Verilog code generation (QiMeng-SALV) by leveraging code segments of functionally correct output signal to optimize RL training. The key insight of QiMeng-SALV is to extract verified signal-aware implementations in partially incorrect modules, so as to enhance the extraction of meaningful functional rewards. Specifically, to verify the functional correctness of signals, some random test inputs are generated to compare signal-level outputs between generated module and reference module in the training data. Additionally, abstract syntax tree (AST) analysis is employed to precisely identify preferred and dispreferred code segments, thus these partially correct signal-level implementation can be utilized to generate meaningful functional rewards. Finally, to enable the model to learn correct signal implementations from erroneous modules, we introduce signal-aware DPO which is based on DPO, a method renowned for its computational efficiency, rapid convergence, and strong performance. Unlike standard DPO that computes probabilities for all module code tokens, signal-aware DPO exclusively calculates token probabilities for the correct signal-level code segments from the preferred and dispreferred samples, thereby preventing noise and interference from incorrect signals during RL training. Crucially, the proposed QiMeng-SALV expands the effective training dataset by incorporating all modules containing any valid signal implementations, regardless of overall module correctness. This represents a fundamental shift from conventional module-level to fine-grained signal-level optimization in RL-based Verilog code generation, addressing the issue of insufficient functional rewards in RL training. To the best of our knowledge, this is the first fine-grained signal-level reinforcement learning algorithm designed for Verilog code generation.
@@ 2.2 RL Training for LLM
In the domain of Verilog generation, current RL-based training approaches face two key challenges: (1) highly rely on structural similarity metrics wang2024veriseek ; wu2024itertl that cannot properly evaluate functional equivalence between different implementations, and (2) the scarcity of high-quality SFT data limits the initial model capability needed for effective RL optimization. While recent work has explored various reward formulations chen2021evaluatingeg2 ; le2022coderl ; wang2024veriseek and feedback mechanisms liu2023rltf ; dou2024stepcoder ; cui2024origen , these approaches often introduce evaluation biases, limiting their reliability in assessing functional correctness. VeriPrefer wang2025insights mitigates such biases by employing module-level functional rewards during reinforcement learning. Our approach differs fundamentally: we adopt signal-level functional rewards, allowing partially correct modules to still provide valuable feedback during training when they contain correctly implemented signals. Consequently, our method delivers denser and more informative reward signals, leading to superior reinforcement learning performance.
@@ 3.2 Signal-aware Verification
First, we use Yosys wolf2013yosys to analyze the reference module head, automatically generating $N$ sets of input signals at equal time intervals. Second, the input signals are simultaneously applied to the reference module $y$ and generated modules $E=\{y_{1},y_{2},...,y_{k}\}$. By comparing the output signals produced by the generated modules with the corresponding output signals from the reference module, we can verify the correctness of generated modules, and obtain the pairs $\{(y_{1},c_{1}),(y_{2},c_{2}),..,(y_{k},c_{k})\}$, where $c_{i}$ is the correct output signals for generated module $y_{i}$. Notably, only output signals exhibiting complete matching across all $N$ sets of input signals are verified as correct, while any discrepancies result in classification as erroneous signals. Third, to select preference samples suitable for DPO training, we choose sample pairs that there exists a set of contrast signals that is correct in the preferred sample but incorrect in the dispreferred sample. This can be formally expressed as:
@@ 3.3 Signal-aware Code Extraction
Based on the above analysis, in the signal-aware code extraction stage, for given target signals, we parse the Abstract Syntax Tree (AST) to establish dependency relationships between signals, and then extract the complete code segment corresponding to the target signals, as depicted in Fig.2 (c). Specifically, given the module code of the target signal, we use Yosys wolf2013yosys to parse the module code and obtain the AST. Subsequently, we analyze the AST to derive the dependency relationships between all the output signals and intermediate signals defined in the module code, forming a signal topology graph. By analyzing this topology graph and performing backward traversal from the target output signal leaf nodes, we obtain the dependent signals of the target signal. Based on the locations of these dependent signals in the AST, we retain their corresponding code segments, thereby obtaining the complete code segment of the target signal. To facilitate the model’s learning by using code segment of contrast signal during DPO training, given the contrast signals $c$, we isolates the relevant code segment $S_{w}^{c}$ from the preferred sample $y_{w}$. Meanwhile, to enable comparison with the preferred sample, we also extract the code segment $S_{l}^{c}$ of contrast signal $c$ from the dispreferred sample $y_{l}$.
@@ 3.4 Signal-aware DPO Training
where $y_{w}$ and $y_{l}$ represents the preferred and dispreferred sample respectively. $S_{w}^{c}$ denotes code segments related to contrast signals $c$ in preferred samples, while $S_{l}^{c}$ represents the same signal related segments in dispreferred samples. $\pi_{\theta}$ is the policy model to be optimized and $\pi_{\text{ref}}$ is the reference model used for regularizing $\pi_{\theta}$ with Kullback-Leibler divergence and $\beta$ is a constant to control the degree of regularization. Through the objective function in Eq. (2), signal-aware DPO shift DPO’s learning focus from entire modules to individual signals, refining the learning granularity and increasing the quantity of functional correctness rewards in DPO, thereby enhancing optimization effectiveness.
@@ 4.1 Experiment Setup
Datasets  Our training dataset is sourced from CodeV zhao2024codev , consisting of 165k samples obtained by crawling all publicly available Verilog module code on GitHub. However, we observed that a subset of these modules are not syntactically valid, prompting us to filter out such cases and retain a cleaned dataset of 135k samples. The 135k dataset is initially employed to fine-tune the general-purpose code generation model Qwen2.5 Coder Instruct hui2024qwen2 , yielding a naive code generator. For every prompt in the 135k dataset, this generator produces 5 candidate module codes, while the corresponding module codes in the dataset serve as reference codes to validate signal correctness during Signal-aware Verification.
Training Settings   In training, we perform full-parameter fine-tuning for 2 epochs during the SFT, followed by about 1 epoch (7000 steps) of LoRA-basedhu2022lora fine-tuning in the Signal-aware DPO. Optimization is carried out using the Adam kingma2014adam optimizer with a cosine annealing learning rate schedule, where the initial learning rates are set to 1.0e-5 for SFT and 5.0e-6 for Signal-aware DPO. The training configuration employs a global batch size of 64 and a maximum sequence length of 2048 tokens.
@@
Baseline Methods  In our experimental evaluation, we conduct a comprehensive comparison between our proposed method and several baseline approaches, categorized into three groups: 1) General-purpose Foundation Models: GPT-3.5, GPT-4o, GPT-4 achiam2023gpt4 , and DeepSeek-v3 liu2024deepseekv3 , which demonstrate broad capabilities across diverse domains. 2) General Code Models: CodeQwen1.5 qwen1.5 , Qwen2.5 Coder Instruct hui2024qwen2 and Deepseek Coder guo2024deepseekcoder , possessing strong coding proficiency but lacking Verilog-specific optimization. 3) Domain-Specialized Verilog Models: Including RTL Coder liu2024rtlcoder , CodeV zhao2024codev , Origen cui2024origen , VeriSeek wang2024veriseek and VeriPrefer wang2025insights which employ module-level rewards in reinforcement learning.
@@ 4.3 Ablation Study
In this section, we present a systematic ablation study to evaluate our proposed QiMeng-SALV. Given that RTLLM contains more complex problems than VerilogEval, making it more suitable for revealing nuanced performance differences among models, we focus our ablation analysis on the RTLLM benchmark. Our investigation spans two key dimensions: 1). configuration variations of Signal-aware DPO training, and 2). different training stage implementations of QiMeng-SALV.
Ablation of configuration variations of signal-aware DPO Unlike standard DPO, which is limited to processing complete correct modules and must consider all response tokens, our signal-aware DPO can not only handle complete correct modules but also extract functional reward in partial correct modules containing correct signal implementations. Therefore, we design three different options to evaluate the efficiency of our method: 1) Complete correct datasets (263k): all of preferred sample in DPO training data are complete correct modules. 2) Partial correct datasets (110k): all of preferred sample in DPO training data are partial correct modules containing both correct and incorrect signals. 3) Filter incorrect signals: determines whether erroneous signals in preferred samples participate in DPO computation. As evidenced by the results in Table 3, standard DPO trained exclusively on complete correct datasets achieves superior performance compared to training on partially correct datasets. Interestingly, combining both complete and partially correct datasets during training leads to degraded performance relative to using only complete correct datasets. This degradation likely stems from the presence of incorrect signal implementations in partially correct module, which introduce noise and hinder the model’s ability to discern meaningful reward signals. By contrast, our proposed signal-aware DPO method, which actively filters out incorrect signal implementations, substantially improves performance in both scenarios: training on partially correct datasets alone and training on mixed datasets. These findings highlight two key advantages of signal-aware DPO: (1) its ability to leverage a broader spectrum of training data, and (2) its robustness against noisy signal implementations through effective filtering.
@@
Scaling on Signal-aware DPO training datasets  To examine the influence of training dataset size on Signal-aware DPO’s efficacy, we conducted experiments by training models with preference datasets of different scales. Figure 3 (b) reveals a consistent upward trend in pass@1, pass@5, and pass@10 metrics as the dataset size grows. These results indicate that incorporating more meaningful functional rewards during reinforcement learning training significantly contributes to improved model performance.